OpenStudy (anonymous):

Write a function rule for the area A of a triangle whose base, b, is 2 cm less than seven times the height, h. What is the area of the triangle when the height is 14 cm?

OpenStudy (campbell_st):

let the height be h... and then the base is 7h - 2

OpenStudy (anonymous):

what would this equal out to be

OpenStudy (campbell_st):

so the area is A = 1/2 b x h using the information above it can be written as A = 1/2 x (7h - 2) x h now just substitute h = 14 and evaluate